OpenStudy (anonymous):

solve the system y=x+2 and y=-x+0

OpenStudy (anonymous):

y = x +2 x = y - 2 y = -(y - 2) + 0 y = -y + 2 2y = 2 2y/2 = 2/2 y = 1 x = (1) - 2 x = -1 (-1,1)

OpenStudy (anonymous):

5*x=20 x=?

OpenStudy (anonymous):

thank you so mutch

OpenStudy (anonymous):

ufail

OpenStudy (anonymous):

you're welcome.